OpenStudy (anonymous):

x2 = 16, therefore x = 4. Is this a valid conclusion? If no, give a counterexample

Directrix (directrix):

False. What other value squared = 16 other than 4? It will be your counterexample.

Directrix (directrix):

Postitive thinking is good for mental health but negative thinking in mathematics sometimes pays off if you think about negative numbers. (strong hint)

OpenStudy (inkyvoyd):

imaginary thinking works best for me.

Directrix (directrix):

She promised to do this but left. The counterexample is -4. x2 = 16 x^2 - 16=0 (x + 4)(x -4) = 0 x = -4 or x = -6. Every number has two square roots.